Part A Five different substances are given to you to be dissolved in water. Which substances...

Part A

Five different substances are given to you to be dissolved in water. Which substances are most likely to undergo dissolution in water?

Check all that apply.

sodium fluoride, NaF
propanol, CH3CH2CH2OH
hexane, C6H14
potassium iodide, KI
heptane, C7H16

Part B

Gastric acid pH can range from 1 to 4, and most of the acid is HCl.

For a sample of stomach acid that is 8.22×10−3 M in HCl, how many moles of HCl are in 19.7 mL of the stomach acid?

Part C

Calculate the molarity of a solution made by putting 55.8 g of NaNO3 into a beaker and diluting to 2.50 L.

Express your answer to three significant figures including the appropriate units.
Answer #1

Part A

All ionic or polar compound will dissolve easily in the water

NaF , CH3CH2CH2OH , KI

Part B

Moles of HCl = molarity * volume ( in litre )

= 8.22*10^-3 * ( 19.7*10^-3)

= 1.62*10^-4 moles

Part C

Moles = mass / molar mass

= 55.8 / 85

= 0.6565

Molarity = moles / volume ( in litres )

= 0.6565 / 2.5

= 0.2626 or 0.263 M

If you have any query please comment

If you satisfied with the solution please like it thankxx